If y(t)=c*e^(-2t)+t+1 is general solution to y'+p(t)*y=g(t), then p(t) and g(t)=?
1
Expert's answer
2020-12-15T03:48:15-0500

Let "y(t)=Ce^{-2t}+t+1" be the general solution of "y'+p(t)y=g(t)".

Then "y'(t)=-2Ce^{-2t}+1". Taking into account that "y'(t)+2y(t)=-2Ce^{-2t}+1+2(Ce^{-2t}+t+1)=2t+3", we conclude that

"p(t)=2" and "g(t)=2t+3."